OpenStudy (anonymous):

Hello! Ok so the question is "Predict what kind of solution would be formed when each of the following oxides reacts with water:" one of the oxides are bromine monoxide, how do I write the equation for this, and is there a formula or something I can remember to help me when a question like this comes up again? Thanks!

OpenStudy (photon336):

well it would have to be \[Br_{2}O\] Because oxygen needs 2 electrons to complete its octet. and bromine needs 1. so it this must be the formula written above. \[Br_{2}O + H_{2}O \rightarrow ~?\]

OpenStudy (photon336):

@Fay-Fay What do you think the product might be?

OpenStudy (anonymous):

I believe it could be Br2O + H2O = BrOH ? However I'm not too sure because before I thought it was Br2O + H2O = HBrO + BrO I'm really not sure.. @Photon336

OpenStudy (photon336):



OpenStudy (photon336):

You notice something right? bromine needs 1 electron to fill its octet while oxygen needs two electrons. so oxygen makes a bond with bromine and a bond with hydrogen

OpenStudy (photon336):

\[Br_{2}O + H_{2}O \rightarrow ~2BrOH\] we put the 2 in front of BrOH so that we can balance the equation.

OpenStudy (anonymous):

I see :) now, what got me confused is the reaction of nitrogen with water, nitrogen reacting with water creates the equation of 3NO2 + H2O = 2HNO3 + NO why is that equation different from the reaction of water with bromine monoxide?

OpenStudy (anonymous):

@phonon336